Grade 12th passGeneral Physics

Dimensional formula for damping constant please reply

Velayutham
7 Years agoGrade 12th pass

2 Answers

Arun
7 Years ago
For the metric system the units are newtons (N) for the force, kilograms (kg) for mass, meters (m) for length and s seconds for time, therefore the speed is given by m/s, the acceleration is m/s2, the spring constant has units N/m, and the damping coeffi- cient is measured in Ns/m = kg/s
Khimraj
7 Years ago

The above is unit less. How? well damping is always force/speed thus c=[Nsm−1], and stiffness is force/distance k=[Nm−1], and of course a newton is [N]=[kgms−2][N]=[kgms−2]

Combine them all to make

\frac{N/(m/s)}{\sqrt{Kg N/m}} =1